OpenStudy (anonymous):

How to solve for x: (2x+18)/6=(2x+16)/2

OpenStudy (anonymous):

x = -15/2

OpenStudy (chaise):

\[\frac{a}{b}=\frac{c}{d}\] becomes: \[\frac{d}{b}=\frac{c}{a}\] Does this help?

OpenStudy (anonymous):

Not really, the answer book says that you need to multiply the 2x+16 by 3 so: 3(2x+16)

OpenStudy (chaise):

Yes, you could do that, so you need to multiply both sides by six. (2x+18)/6*6=(2x+16)/2*6 2x+18=3(2x+16) Does this help you solve for x?

OpenStudy (anonymous):

Oh I think because u multiply both sides by 6 so 1/6 X 6=1 and 1/2 x 6 =3

OpenStudy (chaise):

Yes! Correct.

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

Okay thanks alot.What I did wrong was doing 6(2x+18) and 6(2x+16)

OpenStudy (chaise):

Don't forget you need to still divide the equation by 6 and 2 respectively, so: 6(2x+18)/6 and 6(2x+16)/2

OpenStudy (anonymous):

Oh now It's clear

OpenStudy (anonymous):

\[\frac{ 1 }{ 6 }(2x+18)=\frac{ 1 }{ 2 }(2x+16)\]

OpenStudy (anonymous):

\[x=\frac{ -15 }{ 2 }\]